# Exploring Multilingual Human Value Concepts in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

### v1

@@ Appendix E Complete Results of Cross-Lingual Concept Consistency and Extra Discussions
Results across Model Layers
Effect of Model Size
Regarding model size, despite larger models being able to capture more explicit concepts of human values (as shown in Figure 2), the increase in model size does not steadily enhance cross-lingual concept consistency.
@@
Appendix F Complete Results of Cross-Lingual Concept Transferability and Extra Discussions
Complete Results
@@ Appendix F Complete Results of Cross-Lingual Concept Transferability and Extra Discussions
Effect of Multilinguality
@@
Appendix G Hyperparameter Search and Control Effectiveness Evaluation in Experiments of the Cross-Lingual Value Alignment Control
@@
Figure 7: Cross-lingual similarity of concept vectors of all models across all language pairs, averaged across all human values.
Figure 8: Cross-lingual concept transferability of all models across all language pairs, averaged across all human values.

### v2

@@ G.3 Complete Results
G.4 Effect of Model Size
Despite larger models being able to capture more explicit concepts of human values (as shown in Figure 14), the increase in model size does not steadily enhance cross-lingual concept consistency, as shown in Figure 6.
[…]
LLaMA2-chat-7B 27.3% 22.7% 3.0% 47.0%
Qwen-chat-7B 30.3% 19.7% 7.6% 42.4%
BLOOMZ-7B1 34.1% 15.9% 16.7% 33.3%
Table 7: Proportion of cases in which the concept recognition performance of language A either surpasses or underperforms language B, and whether the transfer from language A to language B is effective or not. $\geq$” and “$<$” denote superiority and inferiority respectively, and $\surd$” and “$\times$” represent successful and unsuccessful transfer.
@@ G.4 Effect of Model Size
Appendix H More Results of Cross-Lingual Concept Transferability
H.1 Transferability Beyond Language Performance
While the setting described in Section 3.4 may introduce bias of initial performance variations across languages, potentially leading to mono-directional transfer from high-performing languages to low-performing ones, our findings suggest that transferability is not solely determined by language performance, as detailed below.
[…]
In summary, although evaluating transferability based solely on changes in accuracy may pose limitations, the phenomenon that transfer is not solely determined by language performance indicates that this remains an open question. We plan to develop more robust and unbiased methodologies to further investigate cross-lingual transfer in our future research.
H.2 Complete Results
@@
H.3 Effect of Multilinguality and Model Size
@@
Appendix I Hyperparameter Search and Control Effectiveness Evaluation in Experiments of the Cross-Lingual Value Alignment Control
@@
Figure 7: Cross-lingual similarity of concept vectors of all models across all language pairs, averaged across all value concepts.
Figure 8: Cross-lingual concept transferability of all models across all language pairs, averaged across all value concepts.